Find an equation of the line that is parallel to y=6x-12 and passes through (-4,1) .
Answer by khurshidi(7)

Find an equation of the line that is parallel to y=6x-12 and passes through (-4,1)?
Solution:
Given line
y = 6x -12
or
6x - y -12 = 0
therefore
Slope of the given line
m1 = - coefficient of x/coefficient of y
coefficient of x = 6, coefficient of y = -1
so
m1 = 6
since the required line is parallel to the given line
therefore
m1 = m2 (m2 = the slope of required line)
i.e. m2 = 6
since the line passes through (-4, 1)= ( x1, y1)
therefore
y - y1 = m2(x - x1)
or
y - 1 = 6(x + 4)
or
y - 1 = 6x + 24
or
6x - y + 25 = 0
This is the required equation parallel to the given equation and passes through (-4, 1).